# NOLO: Navigate Only Look Once

### v1

@@ D.3 Cross Evaluation Results
FloorPlan_Train10_5 60.00$\pm$ 0.96 26.63$\pm$ 1.37 303.93$\pm$ 7.01 2.30$\pm$ 0.13
FloorPlan_Train11_5 64.89$\pm$ 3.50 31.85$\pm$ 3.64 286.70$\pm$ 20.00 2.09$\pm$ 0.22
FloorPlan_Train12_5 52.16$\pm$ 5.79 26.13$\pm$ 3.37 326.15$\pm$ 15.30 2.36$\pm$ 0.24
FloorPlan_Train1_5 50.00$\pm$ 6.24 29.40$\pm$ 1.07 320.91$\pm$ 10.47 2.46$\pm$ 0.04
FloorPlan_Train2_5 47.94$\pm$ 4.56 20.57$\pm$ 4.91 339.62$\pm$ 26.92 2.42$\pm$ 0.18
FloorPlan_Train3_5 39.22$\pm$ 3.09 18.48$\pm$ 0.60 377.95$\pm$ 6.26 3.30$\pm$ 0.22
FloorPlan_Train4_5 64.24$\pm$ 1.87 30.56$\pm$ 0.79 279.33$\pm$ 3.76 2.13$\pm$ 0.13
FloorPlan_Train5_5 61.52$\pm$ 0.86 27.48$\pm$ 2.16 305.47$\pm$ 7.33 2.06$\pm$ 0.11
FloorPlan_Train6_5 60.56$\pm$ 2.83 31.06$\pm$ 4.18 285.57$\pm$ 25.83 2.30$\pm$ 0.16
FloorPlan_Train7_5 50.83$\pm$ 0.59 20.54$\pm$ 1.48 334.78$\pm$ 0.93 2.30$\pm$ 0.17
FloorPlan_Train8_5 54.58$\pm$ 4.12 28.68$\pm$ 1.69 297.57$\pm$ 10.62 2.41$\pm$ 0.02
FloorPlan_Train9_5 38.67$\pm$ 6.60 22.46$\pm$ 4.39 364.06$\pm$ 18.55 2.53$\pm$ 0.15
@@
FloorPlan_Val1_1 54.00$\pm$ 4.32 28.45$\pm$ 3.73 316.52$\pm$ 19.28 2.34$\pm$ 0.15
FloorPlan_Val1_2 48.48$\pm$ 4.09 23.46$\pm$ 3.96 338.57$\pm$ 19.55 2.61$\pm$ 0.12
FloorPlan_Val1_3 74.04$\pm$ 4.06 39.60$\pm$ 2.33 235.18$\pm$ 8.91 1.71$\pm$ 0.09
FloorPlan_Val1_4 44.07$\pm$ 5.83 23.73$\pm$ 1.69 350.40$\pm$ 17.36 2.65$\pm$ 0.12
FloorPlan_Val1_5 49.44$\pm$ 4.10 27.90$\pm$ 4.66 324.46$\pm$ 26.83 2.68$\pm$ 0.17
FloorPlan_Val2_1 32.73$\pm$ 3.71 14.01$\pm$ 1.79 404.18$\pm$ 12.65 3.19$\pm$ 0.10
FloorPlan_Val2_2 48.89$\pm$ 0.91 24.29$\pm$ 3.87 338.47$\pm$ 12.65 2.96$\pm$ 0.19
FloorPlan_Val2_3 67.18$\pm$ 1.92 35.17$\pm$ 2.13 278.73$\pm$ 10.96 1.63$\pm$ 0.08
FloorPlan_Val2_4 52.69$\pm$ 0.38 25.33$\pm$ 0.32 340.45$\pm$ 7.34 2.33$\pm$ 0.03
FloorPlan_Val2_5 44.35$\pm$ 2.61 22.94$\pm$ 1.82 353.16$\pm$ 22.72 2.94$\pm$ 0.11
FloorPlan_Val3_1 59.05$\pm$ 2.86 29.58$\pm$ 2.96 279.42$\pm$ 18.76 2.24$\pm$ 0.23
FloorPlan_Val3_2 56.39$\pm$ 1.71 22.96$\pm$ 1.40 310.36$\pm$ 17.14 2.40$\pm$ 0.08
FloorPlan_Val3_3 59.51$\pm$ 4.58 26.33$\pm$ 3.53 297.91$\pm$ 19.58 2.22$\pm$ 0.14
FloorPlan_Val3_4 41.59$\pm$ 1.19 15.47$\pm$ 1.16 360.06$\pm$ 9.11 2.95$\pm$ 0.01
FloorPlan_Val3_5 60.00$\pm$ 2.36 27.78$\pm$ 2.98 298.36$\pm$ 11.92 2.15$\pm$ 0.15
@@
Table 19: NOLO Performance in Habitat after training in RoboTHOR
@@
Appendix E Social Impact and Future Work

### v2

@@ 9.2 Habitat Evaluation Performance
2azQ1b91cZZ 37.18$\pm$ 4.28 22.84$\pm$ 3.63 359.71$\pm$ 17.26 4.38$\pm$ 0.69
8194nk5LbLH 36.97$\pm$ 1.55 19.47$\pm$ 0.99 366.57$\pm$ 4.66 2.97$\pm$ 0.34
EU6Fwq7SyZv 34.00$\pm$ 2.16 18.48$\pm$ 0.86 378.66$\pm$ 2.54 3.18$\pm$ 0.31
QUCTc6BB5sX 40.33$\pm$ 1.25 23.52$\pm$ 0.77 355.46$\pm$ 1.65 4.83$\pm$ 0.32
SN83YJsR3w2 15.00$\pm$ 7.07 6.92$\pm$ 2.91 455.47$\pm$ 15.62 4.56$\pm$ 1.16
TbHJrupSAjP 20.95$\pm$ 5.99 6.05$\pm$ 1.97 437.43$\pm$ 14.11 4.53$\pm$ 0.51
UwV83HsGsw3 26.67$\pm$ 1.35 17.00$\pm$ 1.76 401.74$\pm$ 3.26 4.77$\pm$ 0.41
Vt2qJdWjCF2 19.17$\pm$ 5.62 8.34$\pm$ 0.78 437.86$\pm$ 13.29 11.49$\pm$ 0.85
VzqfbhrpDEA 25.19$\pm$ 1.05 15.50$\pm$ 2.16 398.60$\pm$ 9.64 5.91$\pm$ 1.16
X7HyMhZNoso 59.05$\pm$ 2.94 30.08$\pm$ 0.75 290.51$\pm$ 7.25 2.00$\pm$ 0.27
Z6MFQCViBuw 38.00$\pm$ 5.66 15.43$\pm$ 1.57 375.89$\pm$ 16.03 5.19$\pm$ 0.22
gTV8FGcVJC9 41.33$\pm$ 1.70 25.04$\pm$ 1.93 335.27$\pm$ 13.56 4.63$\pm$ 0.73
gYvKGZ5eRqb 11.67$\pm$ 2.36 3.51$\pm$ 1.45 474.47$\pm$ 10.70 10.91$\pm$ 0.46
oLBMNvg9in8 44.33$\pm$ 2.36 25.58$\pm$ 1.67 332.98$\pm$ 8.66 1.85$\pm$ 0.12
pLe4wQe7qrG 52.22$\pm$ 3.14 26.99$\pm$ 1.25 316.10$\pm$ 10.12 1.54$\pm$ 0.48
pa4otMbVnkk 14.44$\pm$ 5.67 7.67$\pm$ 3.04 447.38$\pm$ 20.68 7.15$\pm$ 1.53
rqfALeAoiTq 36.94$\pm$ 4.63 18.40$\pm$ 1.58 370.27$\pm$ 11.73 2.45$\pm$ 0.07
x8F5xyUWy9e 46.97$\pm$ 4.71 21.66$\pm$ 1.96 335.51$\pm$ 11.35 2.43$\pm$ 0.57
yqstnuAEVhm 38.61$\pm$ 6.17 17.60$\pm$ 3.23 372.49$\pm$ 22.68 2.89$\pm$ 0.55
zsNo4HB9uLZ 32.50$\pm$ 3.54 18.14$\pm$ 0.90 382.85$\pm$ 11.49 3.92$\pm$ 0.54
Average 33.58 17.41 381.26 4.58
Table 23: NOLO-T performance in Habitat testing scenes.
[…]
2azQ1b91cZZ 41.28$\pm$ 2.97 25.50$\pm$ 2.75 344.03$\pm$ 15.14 4.45$\pm$ 0.08
8194nk5LbLH 39.70$\pm$ 2.61 19.67$\pm$ 0.74 364.86$\pm$ 3.70 3.04$\pm$ 0.27
EU6Fwq7SyZv 41.00$\pm$ 3.27 21.33$\pm$ 2.90 357.03$\pm$ 11.76 3.37$\pm$ 0.17
QUCTc6BB5sX 41.50$\pm$ 4.50 21.52$\pm$ 2.39 357.77$\pm$ 11.40 5.51$\pm$ 0.73
SN83YJsR3w2 21.67$\pm$ 11.79 7.83$\pm$ 4.20 432.92$\pm$ 32.98 5.79$\pm$ 1.89
TbHJrupSAjP 29.52$\pm$ 0.67 11.53$\pm$ 0.91 400.50$\pm$ 5.36 4.45$\pm$ 0.39
UwV83HsGsw3 28.10$\pm$ 0.67 15.16$\pm$ 2.66 398.37$\pm$ 9.11 4.95$\pm$ 0.19
Vt2qJdWjCF2 18.33$\pm$ 3.12 7.74$\pm$ 1.46 443.81$\pm$ 10.96 11.66$\pm$ 0.67
VzqfbhrpDEA 30.00$\pm$ 0.91 16.90$\pm$ 0.92 387.73$\pm$ 5.64 7.28$\pm$ 0.43
X7HyMhZNoso 63.33$\pm$ 0.89 32.24$\pm$ 0.69 270.01$\pm$ 2.15 2.31$\pm$ 0.39
Z6MFQCViBuw 40.67$\pm$ 3.40 15.79$\pm$ 3.40 380.14$\pm$ 11.92 4.87$\pm$ 0.44
gTV8FGcVJC9 46.50$\pm$ 1.50 28.97$\pm$ 1.96 313.77$\pm$ 0.72 5.54$\pm$ 0.94
gYvKGZ5eRqb 10.00 2.73$\pm$ 0.50 471.45$\pm$ 3.05 9.34$\pm$ 1.00
oLBMNvg9in8 47.00$\pm$ 1.00 24.69$\pm$ 1.35 332.24$\pm$ 4.97 1.92$\pm$ 0.16
pLe4wQe7qrG 65.00$\pm$ 1.67 28.04$\pm$ 5.24 290.87$\pm$ 27.90 0.58$\pm$ 0.09
pa4otMbVnkk 33.33$\pm$ 3.33 18.79$\pm$ 0.48 365.55$\pm$ 13.95 7.17$\pm$ 0.66
rqfALeAoiTq 43.89$\pm$ 2.19 22.01$\pm$ 1.75 348.34$\pm$ 8.51 2.55$\pm$ 0.29
x8F5xyUWy9e 52.42$\pm$ 0.43 22.44$\pm$ 0.94 328.36$\pm$ 2.28 2.13$\pm$ 0.02
yqstnuAEVhm 38.61$\pm$ 3.14 17.06$\pm$ 2.82 372.31$\pm$ 14.27 2.83$\pm$ 0.13
zsNo4HB9uLZ 26.67$\pm$ 1.18 16.10$\pm$ 1.78 400.21$\pm$ 5.77 4.74$\pm$ 0.08
Average 37.93 18.80 368.01 4.72
Table 24: NOLO(M) performance in Habitat testing scenes.
[…]
2azQ1b91cZZ 35.77$\pm$ 0.38 21.70$\pm$ 0.98 367.77$\pm$ 0.44 5.26$\pm$ 0.06
8194nk5LbLH 38.18$\pm$ 0.91 19.20$\pm$ 0.74 365.92$\pm$ 3.05 3.05$\pm$ 0.18
EU6Fwq7SyZv 40.50$\pm$ 4.50 20.97$\pm$ 1.42 360.27$\pm$ 14.82 3.04$\pm$ 0.49
QUCTc6BB5sX 40.00 22.11$\pm$ 1.36 354.65$\pm$ 1.51 4.64$\pm$ 0.25
SN83YJsR3w2 32.50$\pm$ 7.50 10.28$\pm$ 2.55 414.92$\pm$ 23.22 4.92$\pm$ 1.20
TbHJrupSAjP 24.29$\pm$ 1.43 7.48$\pm$ 0.50 427.30$\pm$ 6.51 4.63$\pm$ 0.07
UwV83HsGsw3 30.71$\pm$ 2.14 17.89$\pm$ 3.05 390.06$\pm$ 15.81 4.03$\pm$ 0.21
Vt2qJdWjCF2 13.75$\pm$ 1.25 6.65$\pm$ 0.62 454.84$\pm$ 0.97 12.59$\pm$ 0.18
VzqfbhrpDEA 28.33$\pm$ 3.89 14.89$\pm$ 2.38 402.32$\pm$ 9.62 7.08$\pm$ 0.47
X7HyMhZNoso 64.64$\pm$ 1.79 30.95$\pm$ 0.11 276.48$\pm$ 5.50 1.94$\pm$ 0.15
Z6MFQCViBuw 37.00$\pm$ 7.00 16.86$\pm$ 3.37 374.19$\pm$ 22.49 4.90$\pm$ 0.33
gTV8FGcVJC9 50.50$\pm$ 7.50 28.67$\pm$ 1.23 306.25$\pm$ 9.88 4.54$\pm$ 0.72
gYvKGZ5eRqb 12.50$\pm$ 2.50 3.79$\pm$ 1.73 475.62$\pm$ 14.77 9.55$\pm$ 1.20
oLBMNvg9in8 48.00$\pm$ 2.00 28.67$\pm$ 0.73 319.90$\pm$ 6.13 1.61$\pm$ 0.16
pLe4wQe7qrG 53.33$\pm$ 3.33 26.82$\pm$ 1.01 311.75$\pm$ 0.38 1.26$\pm$ 0.19
pa4otMbVnkk 20.00 8.13$\pm$ 0.40 431.58$\pm$ 1.42 8.87$\pm$ 0.33
rqfALeAoiTq 45.42$\pm$ 2.08 20.42$\pm$ 2.06 352.25$\pm$ 7.54 2.19$\pm$ 0.23
x8F5xyUWy9e 49.55$\pm$ 4.09 22.50$\pm$ 1.80 331.82$\pm$ 12.30 2.14$\pm$ 0.47
yqstnuAEVhm 42.08$\pm$ 4.58 18.98$\pm$ 3.12 360.02$\pm$ 16.95 2.89$\pm$ 0.05
zsNo4HB9uLZ 27.50 16.36$\pm$ 1.47 400.51$\pm$ 1.49 4.76$\pm$ 0.10
Average 36.73 18.17 373.92 4.69
9.3 Cross Evaluation Results
The performances of H2R and R2H are reported in Tabs. 25 and 26 respectively.
Table 25: NOLO performance in RoboTHOR after training in Habitat. “T” denotes FloorPlan_Train and “V” denotes FloorPlan_Val.
[…]
T10_5 60.00$\pm$ 0.96 26.63$\pm$ 1.37 303.93$\pm$ 7.01 2.30$\pm$ 0.13
T11_5 64.89$\pm$ 3.50 31.85$\pm$ 3.64 286.70$\pm$ 20.00 2.09$\pm$ 0.22
T12_5 52.16$\pm$ 5.79 26.13$\pm$ 3.37 326.15$\pm$ 15.30 2.36$\pm$ 0.24
T1_5 50.00$\pm$ 6.24 29.40$\pm$ 1.07 320.91$\pm$ 10.47 2.46$\pm$ 0.04
T2_5 47.94$\pm$ 4.56 20.57$\pm$ 4.91 339.62$\pm$ 26.92 2.42$\pm$ 0.18
T3_5 39.22$\pm$ 3.09 18.48$\pm$ 0.60 377.95$\pm$ 6.26 3.30$\pm$ 0.22
T4_5 64.24$\pm$ 1.87 30.56$\pm$ 0.79 279.33$\pm$ 3.76 2.13$\pm$ 0.13
T5_5 61.52$\pm$ 0.86 27.48$\pm$ 2.16 305.47$\pm$ 7.33 2.06$\pm$ 0.11
T6_5 60.56$\pm$ 2.83 31.06$\pm$ 4.18 285.57$\pm$ 25.83 2.30$\pm$ 0.16
T7_5 50.83$\pm$ 0.59 20.54$\pm$ 1.48 334.78$\pm$ 0.93 2.30$\pm$ 0.17
T8_5 54.58$\pm$ 4.12 28.68$\pm$ 1.69 297.57$\pm$ 10.62 2.41$\pm$ 0.02
T9_5 38.67$\pm$ 6.60 22.46$\pm$ 4.39 364.06$\pm$ 18.55 2.53$\pm$ 0.15
@@ 9.3 Cross Evaluation Results
V1_1 54.00$\pm$ 4.32 28.45$\pm$ 3.73 316.52$\pm$ 19.28 2.34$\pm$ 0.15
V1_2 48.48$\pm$ 4.09 23.46$\pm$ 3.96 338.57$\pm$ 19.55 2.61$\pm$ 0.12
V1_3 74.04$\pm$ 4.06 39.60$\pm$ 2.33 235.18$\pm$ 8.91 1.71$\pm$ 0.09
V1_4 44.07$\pm$ 5.83 23.73$\pm$ 1.69 350.40$\pm$ 17.36 2.65$\pm$ 0.12
V1_5 49.44$\pm$ 4.10 27.90$\pm$ 4.66 324.46$\pm$ 26.83 2.68$\pm$ 0.17
V2_1 32.73$\pm$ 3.71 14.01$\pm$ 1.79 404.18$\pm$ 12.65 3.19$\pm$ 0.10
V2_2 48.89$\pm$ 0.91 24.29$\pm$ 3.87 338.47$\pm$ 12.65 2.96$\pm$ 0.19
V2_3 67.18$\pm$ 1.92 35.17$\pm$ 2.13 278.73$\pm$ 10.96 1.63$\pm$ 0.08
V2_4 52.69$\pm$ 0.38 25.33$\pm$ 0.32 340.45$\pm$ 7.34 2.33$\pm$ 0.03
V2_5 44.35$\pm$ 2.61 22.94$\pm$ 1.82 353.16$\pm$ 22.72 2.94$\pm$ 0.11
V3_1 59.05$\pm$ 2.86 29.58$\pm$ 2.96 279.42$\pm$ 18.76 2.24$\pm$ 0.23
V3_2 56.39$\pm$ 1.71 22.96$\pm$ 1.40 310.36$\pm$ 17.14 2.40$\pm$ 0.08
V3_3 59.51$\pm$ 4.58 26.33$\pm$ 3.53 297.91$\pm$ 19.58 2.22$\pm$ 0.14
V3_4 41.59$\pm$ 1.19 15.47$\pm$ 1.16 360.06$\pm$ 9.11 2.95$\pm$ 0.01
V3_5 60.00$\pm$ 2.36 27.78$\pm$ 2.98 298.36$\pm$ 11.92 2.15$\pm$ 0.15
@@
Table 26: NOLO Performance in Habitat testing scenes after training in RoboTHOR.
@@
10 Social Impact and Future Work